OpenStudy (anonymous):

Write the equation of a circle with center P(0, -3) and radius 8. a.) x2 + (y + 3)2 = 64 b.) x2 + (y - 3)2 = 64 c.) (x - 3)2 + (y + 3)2 = 64 d.) (x - 3)2 + y2 = 64

jimthompson5910 (jim_thompson5910):

Hint: \[\Large (x-h)^2 + (y-k)^2 = r^2\] is the general equation of a circle. The center is (h,k) and the radius is r.

OpenStudy (anonymous):

thank you i got b

jimthompson5910 (jim_thompson5910):

you are close

jimthompson5910 (jim_thompson5910):

but b is incorrect

OpenStudy (anonymous):

ohh um ill try again hang on

jimthompson5910 (jim_thompson5910):

notice in choice b how there is a "y-3" compare y-3 with y-k this tells us that k = 3 and the y coordinate of the center is 3

OpenStudy (anonymous):

is it a?

jimthompson5910 (jim_thompson5910):

but we want the y coordinate of the center to be -3

jimthompson5910 (jim_thompson5910):

it is because y+3 is the same as y - (-3)

OpenStudy (anonymous):

i just realised that its two negatives that make a posotive

jimthompson5910 (jim_thompson5910):

that's one way you can think of it

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome